Use the FOIL method to find the product. (6x^2-4)(x^8-9).
Answer by ankor@dixie-net.com(22740)

Use the FOIL method to find the product. (6x^2-4)(x^8-9).
:
F: 6x^2 * x^8 = 6x^10
O: 6x^2 * -9 = -54x^2
I: -4 * x^8 = -4x^8
L: -4 * -9 = +36
Which in descending order:
6x^10 - 4x^8 - 54x^2 + 36